# The unrealized potential of agroforestry for an emissions-intensive agricultural commodity

### v1

@@ Acknowledgements
Figure S1: Distribution of agro-climatic zones and shade levels for cocoa production in Côte d’Ivoire and Ghana. Different colors represent various climatic types determined from climatic data spanning 1990-2020. Areas in dark grey (mixed) show classification uncertainty, while light grey areas (limitations) are likely unsuitable for cocoa cultivation but with high uncertainty, and white areas are unsuitable for cocoa growth. Black dots indicate cocoa communities visited during fieldwork for ground truth data collection. Sub-panels (a-e) display the shade levels in each agro-climatic zone based on our map of shade-tree cover (Fig. 1).
Figure S2: Differences in climatic variable across five distinct agro-climatic types (1–5). The variables include: (a) isothermality (the ratio of diurnal temperature variation to annual temperature range), (b) temperature seasonality (annual variability in temperature), (c) mean temperature of the wettest quarter, (d) mean temperature of the coldest quarter, (e) precipitation of the driest month, (f) precipitation seasonality (coefficient of variation in monthly precipitation), (g) precipitation of the warmest quarter, (h) precipitation of the coldest quarter, (i) consecutive months with precipitation below 100 mm, (j) potential evapotranspiration (PET) seasonality (annual variability in PET), (k) PET range (difference between maximum and minimum PET), (l) consecutive months with precipitation less than PET, and (m) mean temperature during the wettest semester. The boxplots display the median (central line), interquartile range (box limits), 10th and 90th percentiles (whiskers), and potential outliers (dots). Different colors correspond to the five agro-climatic types identified.
Figure S3: Shade levels in a) cocoa growing districts of Côte d’Ivoire and b) cocoa growing regions of Ghana. The maps display the distribution of administrative districts/regions within cocoa-growing areas, shaded in turquoise. Mean shade levels (%) are shown for each district/region, with error bars representing $\pm$ standard deviations. The bubble sizes correspond to the total cocoa-growing area in each region, ranging from 78,382 to 1,190,000 hectares, with larger bubbles indicate regions with larger cocoa-growing areas. Only administrative regions/districts comprising more than 2% of the total cocoa-growing area in each country were included in the plot.
Figure S4: Area, carbon densities, and carbon stocks of cocoa-growing areas, disturbed forests (dist. forest), and undisturbed forests (forest) across Ghana and Côte d’Ivoire. The values in (a) represent the total area with given land use , for which there are no error estimates. The values in (b) are means $\pm$ standard deviations because we are interested in the variation among locations in carbon density. The values in (c) are total carbon $\pm$ 95 % CIs, because we are interested in the uncertainty in the estimates of total carbon. Because there are more than $3\times 10^{7}$ pixel-level estimates of total carbon, CIs in (c) are too small to be shown. 95 % confidence intervals for total carbon in cocoa: ±$\pm 1.2\times 10$-5 million tonnes, disturbed forest: ±$\pm 4.1\times 10$-6 million tonnes C, and undisturbed forest: ±$\pm 2.5\times 10$-6 million tonnes C. Please note, all values were calculated at the resolution of our AGBD map (50$\times$50 m).
@@
Figure S5: Residual analysis for a) shade tree cover and b) aboveground biomass density (AGBD). a) Residual analysis of shade tree cover intervals (each spanning 5%) from our map, compared against ground truth field measurements. b) Residual analysis of AGBD intervals (each spanning 15 tonnes ha-1) for our map and the European Space Agency’s (ESA) Climate Change Initiative (CCI) Biomass map [31], referenced against GEDI ground truth data. Boxplots show the median, quartiles, and 10th to 90th percentiles (whiskers) of the residuals. Negative residuals indicate that estimates are lower than reference values. Our map generally exhibits lower biases compared to the ESA map, except at very high AGBD values.
Figure S6: Distribution of heights for cocoa trees in ground truth data. Cocoa tree heights were derived from 6,800 manually-labeled polygons across small sections of cocoa monocultures, using drone-derived canopy height maps. These polygons covered 118 hectares and encompassed 230 cocoa farms throughout the study region. An 8-meter height cut-off (indicated by the dashed vertical line) was used to distinguish cocoa trees from shade trees. Notably, 99.7% of all cocoa tree height data points fall below this threshold.
@@
Figure S7: The relationship between shade tree cover and aboveground biomass (AGB) in cocoa growing areas across Ghana and Côte d’Ivoire. The plot is based on pixel-level estimates of shade tree cover and AGB, using a randomly selected 10% subset of the total dataset (approximately 3 million data points). Data were sampled at a 50-meter ground distance and are presented as a hexbin plot. The hexagonal bins represent the density of data points, with dark blue indicating higher density and light green indicating lower density. The solid red line shows the Bayesian regression analysis for shade levels up to 40%, while the red dotted line represents the 95% uncertainty interval based on the posterior predictive distribution’s standard deviation for the same range.

### v2

@@ Acknowledgements
Fig. S1: Residual analysis for a) shade-tree cover and b) aboveground biomass density (AGBD). a) Residual analysis of shade-tree cover intervals (each spanning 5%) from our map, compared against ground truth field measurements. b) Residual analysis of AGBD intervals (each spanning 15 tonnes ha-1) for our map and the European Space Agency’s (ESA) Climate Change Initiative (CCI) Biomass map, referenced against GEDI ground truth data. The boxplots display the median, interquartile range, and whiskers extending from the 10th to 90th percentiles of the residuals. Negative values indicate underestimation relative to the reference. Our map generally exhibits lower biases compared to the ESA map, except at very high AGBD values. Panel a: $n=259$ cocoa farms in the hold-out test set. Panel b: $n=44,220$ GEDI footprints in the hold-out test set. No statistical tests were performed; residuals are shown to assess model accuracy.
@@
Fig. S2: The relationship between shade-tree cover and aboveground biomass (AGB) in cocoa growing areas across Ghana and Côte d’Ivoire. The plot is based on pixel-level estimates of shade-tree cover and AGB, using a randomly selected 10% subset of the total dataset (over 3.4 million data points). Data were sampled at a 50-meter ground distance and are presented as a hexbin plot. The hexagonal bins represent the density of data points, with dark blue indicating higher density and light green indicating lower density. The solid red line shows the posterior median from a Bayesian regression model fit to data with shade cover $\leq$40%, and the red dotted lines show the 95% credible interval based on predictions generated from the posterior distribution. Model details are provided in the Supplementary Methods under Bayesian regression methods for quantifying relationship between cover and biomass.
Fig. S3: Differences in climatic variable across five distinct agro-climatic types (1–5). The bioclimatic variables shown include: (a) isothermality (BIO 3: ratio of diurnal temperature range to annual temperature range), (b) temperature seasonality (BIO 4: standard deviation of temperature ×100), (c) mean temperature of the wettest quarter (BIO 8), (d) mean temperature of the coldest quarter (BIO 11), (e) precipitation of the driest month (BIO 14), (f) precipitation seasonality (BIO 15: coefficient of variation in monthly precipitation), (g) precipitation of the warmest quarter (BIO 18), (h) precipitation of the coldest quarter (BIO 19), (i) number of consecutive months with precipitation below 100 mm (BIO 20), (j) potential evapotranspiration (PET) seasonality (BIO 26: annual variability in PET), (k) PET range (BIO 29: difference between maximum and minimum PET), (l) number of consecutive months with precipitation less than PET (BIO 21), and (m) mean temperature during the wettest semester (BIO 23). The boxplots display the median (central line), interquartile range (box limits), 10th and 90th percentiles (whiskers), and potential outliers (dots). Different colors correspond to the five agro-climatic types identified. $n=936$ cocoa locations across all agro-climatic types. No statistical tests were performed; distributions are shown to visualize climatic differences among clusters identified via Random Forest dissimilarity-based clustering. Bioclimatic variable follow the methods and definitions used in Kamath et al. (2024) [Kamath2024]
Fig. S4: Distribution of heights for cocoa trees in ground truth data. Cocoa tree heights were derived from 6,800 manually-labeled polygons across small sections of cocoa monocultures, using drone-derived canopy height maps. These polygons covered 118 hectares and encompassed 230 cocoa farms throughout the study region. An 8-meter height cut-off (indicated by the dashed vertical line) was used to distinguish cocoa trees from shade trees. Notably, 99.7% of all cocoa tree height data points fall below this threshold. N = 149,913,724 individual pixels within the labeled cocoa polygons. No statistical tests were performed; data are shown to characterize the distribution of cocoa canopy heights